# Coordination Control of Discrete Event Systems under Cyber Attacks

### v2

@@ 1 Introduction
In coordination control, we start with $G$ and $K$. We design two or more local supervisors based on local versions of $G$ and $K$. To do so, we properly select two or more sets of local events to ensure that $K$ satisfies two conditions. The first condition is conditional decomposability, which says that the synchronous product of the projections of $K$ to some sets of local events is equal to $K$. The second condition is the observer property [14], which says that the numbers of states in the projected automata is smaller than the number of states in $G$. The two conditions can always be satisfied by properly augmenting the local alphabets (event sets). When they are satisfied, coordination control is highly effective for supervisory control of DES. It simplifies the design, improves implementation, and ensures scalability while reducing the computational complexity [9, 6, 8]. In this paper, we use coordination control for discrete event system under cyber attacks (CA).
@@ 2.1 Discrete Event Systems and Supervisory Control
We use $\Sigma^{*}$ to denote the set of all (finite) strings over $\Sigma$. The transition function $\delta$ on single event $\sigma$ is extended to strings $s$, that is, $\delta:Q\times\Sigma^{*}\rightarrow Q$ in the usual way [3, 25]. If $\delta(q,s)$ is defined, we denote it by $\delta(q,s)!$. The language generated by $G$ is the set of all strings defined in $G$ from the initial state:
@@
$P(\varepsilon)=$ $\varepsilon,$ $P(s\sigma)=$$\begin{cases}P(s)\sigma&\text{if $\sigma\in\Sigma_{o}$}\\ P(s)&\text{if $\sigma\in\Sigma_{uo}$}\\ \end{cases}$
where $\varepsilon$ is the empty string. The inverse mapping of $P$ is defined as $P^{-1}(w)=\{s\in\Sigma^{*}:P(s)=w\}$. $P$ and $P^{-1}$ are extended to a language in the usual way.
@@ 2.2 Supervisory Control under Cyber Attacks
An attacker can observe observable events. When $\sigma\in\Sigma_{o}^{a}$ in ${tr}=(q,\sigma,q^{\prime})\in\delta^{a}$ occurred in $G$ and is intercepted by the attacker, the attacker can change the event $\sigma$ to any string in an attack language $A_{tr}\subseteq\Sigma^{*}$. Note that, since unobservable events neither help nor confuse the supervisor, whether we require $A_{tr}\subseteq\Sigma^{*}$ or $A_{tr}\subseteq\Sigma_{o}^{*}$ will make no difference.
Hence, if a string $s=\sigma_{1}\sigma_{2}...\sigma_{|s|}\in L(G)$ occurs in $G$, the set of possible strings after sensor attacks, denoted by $\Theta^{a}(s)$, is calculated as follows. Let $q_{k}=\delta(q_{o},\sigma_{1}\cdots\sigma_{k}),tr_{k}=(q_{k-1},\sigma_{k},q_{k}),k=1,2,...,|s|$, then $\Theta^{a}(s)=L_{1}L_{2}...L_{|s|},$ where
$\begin{split}L_{k}=\left\{\begin{array}[]{ll}\{\sigma_{k}\}&\mbox{if }\sigma_{k}\not\in\Sigma_{o}^{a}\\ A_{tr_{k}}&\mbox{if }\sigma_{k}\in\Sigma_{o}^{a}.\end{array}\right.\end{split}$ (1)
In other words, the attacker can replace the event in an attackable transition by any one of the strings in the corresponding attack language. Since $\Theta^{a}(s)$ is a set of strings, $\Theta^{a}$ is a mapping $\Theta^{a}:L(G)\rightarrow 2^{\Sigma^{*}}$.
This sensor attack model is general and includes the following special cases. (1) No attack: if $\sigma\in A_{tr}$ and $\sigma$ is altered to $\sigma$ (no change), then there is no attack. (2) Deletion: if the empty string $\varepsilon\in A_{tr}$ and $\sigma$ is altered to $\varepsilon$, then $\sigma$ is deleted. (3) Replacement: if $\alpha\in A_{tr}$ and $\sigma$ is altered to $\alpha$, then $\sigma$ is replaced by $\alpha$. (4) Insertion: if $\sigma\alpha\in A_{tr}$ (or $\alpha\sigma\in A_{tr}$) and $\sigma$ is altered to $\sigma\alpha$ (or $\alpha\sigma$), then $\alpha$ is inserted. Note that, from practical viewpoint, we may want to assume that $A_{tr}$ is finite, although such an assumption is not needed theoretically.
@@
To model sensor attacks for automaton $G$, we do the following. For each transition ${tr}\in\delta^{a}$, let $A_{tr}$ be marked by an automaton $F_{tr}$, that is, $A_{tr}=L_{m}(F_{tr})$ for some
$F_{tr}=(Q_{tr},\Sigma,\delta_{tr},q_{0,{tr}},Q_{m,{tr}}).$
Note that not all states in $F_{tr}$ are marked. We assume that $F_{tr}$ is trim, that is, all states are accessible from $q_{o,{tr}}$ and co-accessible to $Q_{m,{tr}}$.
Let us first show how to substitute an attackable transition ${tr}=(q,\sigma,q^{\prime})\in\delta^{a}$ by a set of sequences of transitions $((q,\varepsilon,q_{0,tr}),\delta_{tr},(q_{m,tr},\varepsilon,q^{\prime}))$ (with $\delta_{tr}$ being the set of transition sequences defined in $F_{tr}$). We use the intuitive notation $G_{{tr}\rightarrow(q,F_{tr},q^{\prime})}$ to denote this substitution procedure, which is formalized as follows.
$G_{{tr}\rightarrow(q,F_{tr},q^{\prime})}=(Q\cup Q_{tr},\Sigma,\delta_{{tr}\rightarrow(q,F_{tr},q^{\prime})},q_{o},{Q_{m}}),$
where $\delta_{tr\rightarrow(q,F_{tr},q^{\prime})}=(\delta\setminus\{(q,\sigma,q^{\prime})\})\cup\delta_{tr}\cup\{(q,\varepsilon,q_{0,tr})\}$$\cup\{(q_{m,tr},\varepsilon,q^{\prime}):q_{m,tr}\in Q_{m,tr}\}$. In other words, $G_{tr\rightarrow(q,F_{tr},q^{\prime})}$ contains all transitions in $\delta$ and $\delta_{tr}$, except $(q,\sigma,q^{\prime})$, plus the $\varepsilon$-transitions from $q$ to $q_{0,{tr}}$ and from $Q_{m,{tr}}$ to $q^{\prime}$. If $F_{tr}$ has only one marked state, that is, $Q_{m,tr}=\{q_{m,tr}\}$, then we can use a shortcut by merging $q$ and $q^{\prime}$ with $q_{0,tr}$ and $q_{m,tr}$, respectively, without the $\varepsilon$-transitions $(q,\varepsilon,q_{0,tr})$ and $(q_{m,tr},\varepsilon,q^{\prime})$.
@@
where $\tilde{Q}=\cup_{tr\in\delta^{a}}Q_{tr}$ is the set of states added during the substitution and $Q^{a}_{m}=Q$ is the set of marked states. $G^{a}$ models the attacked system by combining $G$ and the ALTER model. Note that $G^{a}$ is a nondeterministic automaton and its transition is a mapping $\delta^{a}:Q^{a}\times\Sigma\rightarrow 2^{Q^{a}}$. From the construction of $G^{a}$, it can be shown [29] that
@@
Sensor attacks can be either “static” or “dynamic”. A sensor attack is static if, for any two transitions ${tr_{1}}=(q_{1},\sigma,q_{1}^{\prime})\in\delta^{a}$ and ${tr_{2}}=(q_{2},\sigma,q_{2}^{\prime})\in\delta^{a}$ with the same event $\sigma$, we have $A_{tr_{1}}=A_{tr_{2}}$. Hence, $A_{tr}$ can also be written as $A_{\sigma}$. A sensor attack is dynamic if the above is not true, that is, $A_{tr_{1}}\not=A_{tr_{2}}$ for some ${tr_{1}}=(q_{1},\sigma,q_{1}^{\prime})\in\delta^{a}$ and ${tr_{2}}=(q_{2},\sigma,q_{2}^{\prime})\in\delta^{a}$ with the same event $\sigma$.
@@
The set of all possible transitions of $M$ is denoted by $\zeta=\{(y,\sigma,y^{\prime}):\zeta(y,\sigma)=y^{\prime}\}$. The set of attackable transitions is denoted by $\zeta^{a}=\{(y,\sigma,y^{\prime})\in\zeta:\sigma\in\Sigma_{o}^{a}\}$. For each transition ${tr}=(y,\sigma,y^{\prime})\in\zeta^{a}$, the attacker can change the event $\sigma$ to any string in a language $A^{M}_{tr}\subseteq\Sigma^{*}$, in a way similar to $A_{tr}$, but for transitions in $M$. Note that, since the attacker observes $\Sigma_{o}$, it knows which state $y\in Y$ it is in. Hence, when it observes an event, it knows which transition the event corresponds to. Therefore, ALTER can be implemented based on $M$.
@@ 3.3 Relation between local and global large languages
Under Assumption 1, the large language generated by $(\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G$ is given by
@@
(by the definition of $(\mbox{by the definition of }L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G))$
@@
(because $(\mbox{because }\sigma\in\Sigma_{1}\setminus\Sigma_{2}\Rightarrow\sigma\not\in\Sigma\setminus\Sigma_{1}\wedge\sigma\in\Sigma\setminus\Sigma_{2})$
@@
[…]
@@
(because $(\mbox{because }\sigma\in\Sigma_{1}\setminus\Sigma_{2})$
@@
[…]
@@
(by the definition of $(\mbox{by the definition of }L_{a}(\mathcal{S}_{1}^{a}/G_{1}))$
@@
(because $(\mbox{because }s\in L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G)\wedge s\sigma\in L(G)\Rightarrow P_{1}(s)\in L_{a}(\mathcal{S}_{1}^{a}/G_{1})\wedge P_{1}(s)\sigma\in L(G_{1}))$
@@
(by the definition of $(\mbox{by the definition of }L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G))$
@@
(because $(\mbox{because }\sigma\in\Sigma_{1}\cap\Sigma_{2}\Rightarrow\sigma\not\in\Sigma\setminus\Sigma_{1}\wedge\sigma\not\in\Sigma\setminus\Sigma_{2})$
@@
(because $(\mbox{because }\sigma\in\Sigma_{1}\cap\Sigma_{2})$
@@
(by the definition of $(\mbox{by the definition of }L_{a}(\mathcal{S}_{i}^{a}/G_{i}))$
@@
(because $(\mbox{because }s\in L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G)\wedge s\sigma\in L(G)\Rightarrow P_{i}(s)\in L_{a}(\mathcal{S}_{i}^{a}/G_{i})\wedge P_{i}(s)\sigma\in L(G_{i}))$
@@ 4 Design of Coordination Control under Cyber Attacks
Under Assumption 1, there exist local supervisors $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that
@@
Step 3. Replace unobservable transitions in $H_{i}^{a}$ by $\varepsilon$-transitions as in Equation (6) and denote the resulting automaton as
@@
[…]
Step 4. To obtain the state estimates, we convert $H_{i,\varepsilon}^{a}$ to a deterministic automaton $H_{i,obs}^{a}$, called CA-observer, as
@@
Step 5. To ensure that the supervised system never enters illegal/unsafe states $Q_{i}\setminus Q_{H,i}$, the set of events that need to be disabled after observing $w\in\Phi_{i}^{a}(L(H_{i}))$ is given by
@@
Consider a discrete event system $G_{i}$ under joint sensor and actuator attacks. For a nonempty closed language $P_{i}(K)\subseteq L(G_{i})$, if $P_{i}(K)$ is CA-controllable (with respect to $L(G_{i})$, $\Sigma_{uc}\cap\Sigma_{i}$, and $\Sigma^{a}_{c}\cap\Sigma_{i}$) and CA-observable (with respect to $L(G_{i})$, $\Sigma_{o}\cap\Sigma_{i}$, and $\Phi_{i}^{a}$), $i=1,2$, then the CA-supervisors $\mathcal{S}_{i,CA}$ defined in Equation (14) achieve $P_{i}(K)$, that is,

### v3

@@ 1 Introduction
In coordination control, we start with $G$ and $K$. We design two or more local supervisors based on local versions of $G$ and $K$. To do so, we properly select two or more sets of local events to ensure that $K$ satisfies two conditions. The first condition is conditional decomposability [9], which says that the synchronous product of the projections of $K$ to some sets of local events is equal to $K$. The second condition is the observer property [14], which says that the numbers of states in the projected automata is smaller than the number of states in $G$. The two conditions can always be satisfied by properly augmenting the local alphabets (event sets). When they are satisfied, coordination control is highly effective for supervisory control of DES. It simplifies the design, improves implementation, and ensures scalability while reducing the computational complexity [9, 6, 8]. In this paper, we use coordination control for discrete event system under cyber attacks (CA).
@@ 2.1 Discrete Event Systems and Supervisory Control
We use $\Sigma^{*}$ to denote the set of all (finite) strings over $\Sigma$, including the empty string $\varepsilon$. The transition function $\delta$ on single event $\sigma$ is extended to strings $s$, that is, $\delta:Q\times\Sigma^{*}\rightarrow Q$ in the usual way [3, 25]. If $\delta(q,s)$ is defined, we denote it by $\delta(q,s)!$. The language generated by $G$ is the set of all strings defined in $G$ from the initial state:
@@
[…]
where $\varepsilon$ is the empty string. The inverse mapping of $P$ is defined as $^{-1}(w)=\{s\in\Sigma^{*}:P(s)=w\}. PP$ and $P^{-1}$ are extended to a language in the usual way.
@@ 2.2 Supervisory Control under Cyber Attacks
An attacker can observe observable events. When $\sigma\in\Sigma_{o}^{a}$ in ${tr}=(q,\sigma,q^{\prime})\in\delta^{a}$ occurred in $G$ and is intercepted by the attacker, the attacker can change the event $\sigma$ to any string in an attack language $A_{\sigma}\subseteq\Sigma^{*}$. Note that, since unobservable events neither help nor confuse the supervisor, whether we require $A_{\sigma}\subseteq\Sigma^{*}$ or $A_{\sigma}\subseteq\Sigma_{o}^{*}$ will make no difference.
Hence, if a string $s=\sigma_{1}\sigma_{2}...\sigma_{|s|}\in L(G)$ occurs in $G$, the set of possible strings after sensor attacks, denoted by $\Theta^{a}(s)$. $\Theta^{a}(s)=A_{\sigma_{1}}A_{\sigma_{2}}...A_{\sigma_{|s|}},$
This sensor attack model is general and includes the following special cases. (1) No attack: if $\sigma\in A_{\sigma}$ and $\sigma$ is altered to $\sigma$ (no change), then there is no attack. That is $A_{\sigma}=\{\sigma,\cdots\}$. (2) Deletion: if the empty string $\varepsilon\in A_{\sigma}$ and $\sigma$ is altered to $\varepsilon$, then $\sigma$ is deleted. That is $A_{\sigma}=\{\varepsilon,\cdots\}$. (3) Replacement: if $\alpha\in A_{\sigma}$ and $\sigma$ is altered to $\alpha$, then $\sigma$ is replaced by $\alpha$. That is $A_{\sigma}=\{\alpha,\cdots\}$. (4) Insertion: if $\sigma\alpha\in A_{\sigma}$ (or $\alpha\sigma\in A_{\sigma}$) and $\sigma$ is altered to $\sigma\alpha$ (or $\alpha\sigma$), then $\alpha$ is inserted. That is $A_{\sigma}=\{\sigma\alpha,\cdots\}$. (5) All-out attack: if $A_{\sigma}=\Sigma^{*}$, then the attacker can change $\sigma$ to any strings.Note that, from practical viewpoint, we may want to assume that $A_{\sigma}$ is finite, although such an assumption is not needed theoretically.
@@
To model sensor attacks for automaton $G$, we do the following. For each attackable event $\sigma\in\Sigma_{o}^{a}$, let $A_{\sigma}$ be marked by an automaton $F_{\sigma}$, that is, $A_{\sigma}=L_{m}(F_{\sigma})$ for some
$F_{\sigma}=(Q_{\sigma},\Sigma,\delta_{\sigma},q_{0,{\sigma}},Q_{m,{\sigma}}).$
Note that not all states in $F_{\sigma}$ are marked. We assume that $F_{\sigma}$ is trim, that is, all states are accessible from $q_{o,{\sigma}}$ and co-accessible to $Q_{m,{\sigma}}$.
Let us first show how to substitute an attackable transition ${tr}=(q,\sigma,q^{\prime})\in\delta^{a}$ by a set of sequences of transitions $((q,\varepsilon,q_{0,\sigma}),\delta_{\sigma},(q_{m,\sigma},\varepsilon,q^{\prime}))$ (with $\delta_{\sigma}$ being the set of transition sequences defined in $F_{\sigma}$). We use the intuitive notation $G_{{(q,\sigma,q^{\prime})}\rightarrow(q,F_{\sigma},q^{\prime})}$ to denote this substitution procedure, which is formalized as follows.
$G_{{(q,\sigma,q^{\prime})}\rightarrow(q,F_{\sigma},q^{\prime})}=(Q\cup Q_{\sigma},\Sigma,\delta_{{(q,\sigma,q^{\prime})}\rightarrow(q,F_{\sigma},q^{\prime})},q_{o},{Q_{m}}),$
where $\delta_{(q,\sigma,q^{\prime})\rightarrow(q,F_{\sigma},q^{\prime})}=(\delta\setminus\{(q,\sigma,q^{\prime})\})\cup\delta_{\sigma}\cup\{(q,\varepsilon,q_{0,\sigma})\}$ $\cup\{(q_{m,\sigma},\varepsilon,q^{\prime}):q_{m,\sigma}\in Q_{m,\sigma}\}$. In other words, $G_{(q,\sigma,q^{\prime})\rightarrow(q,F_{\sigma},q^{\prime})}$ contains all transitions in $\delta$ and $\delta_{\sigma}$, except $(q,\sigma,q^{\prime})$, plus the $\varepsilon$-transitions from $q$ to $q_{0,{\sigma}}$ and from $Q_{m,{\sigma}}$ to $q^{\prime}$. If $F_{\sigma}$ has only one marked state, that is, $Q_{m,\sigma}=\{q_{m,\sigma}\}$, then we can use a shortcut by merging $q$ and $q^{\prime}$ with $q_{0,\sigma}$ and $q_{m,\sigma}$, respectively, without the $\varepsilon$-transitions $(q,\varepsilon,q_{0,\sigma})$ and $(q_{m,\sigma},\varepsilon,q^{\prime})$.
@@
where $\tilde{Q}=\cup_{\sigma\in\delta^{a}}Q_{\sigma}$ is the set of states added during the substitution and $Q^{a}_{m}=Q$ is the set of marked states. $G^{a}$ models the attacked system by combining $G$ and the ALTER model. Note that $G^{a}$ is a nondeterministic automaton and its transition is a mapping $\delta^{a}:Q^{a}\times\Sigma\rightarrow 2^{Q^{a}}$. From the construction of $G^{a}$, it can be shown [29] that
@@
Sensor attacks can be either “static” or “dynamic”. A sensor attack is static if, for any two transitions ${tr_{1}}=(q_{1},\sigma,q_{1}^{\prime})\in\delta^{a}$ and ${tr_{2}}=(q_{2},\sigma,q_{2}^{\prime})\in\delta^{a}$ with the same event $\sigma$, we have $A_{tr_{1}}=A_{tr_{2}}$. Hence, $A_{\sigma}$ can also be written as $A_{\sigma}$. A sensor attack is dynamic if the above is not true, that is, $A_{tr_{1}}\not=A_{tr_{2}}$ for some ${tr_{1}}=(q_{1},\sigma,q_{1}^{\prime})\in\delta^{a}$ and ${tr_{2}}=(q_{2},\sigma,q_{2}^{\prime})\in\delta^{a}$ with the same event $\sigma$.
@@
The set of all possible transitions of $M$ is denoted by $\zeta=\{(y,\sigma,y^{\prime}):\zeta(y,\sigma)=y^{\prime}\}$. The set of attackable transitions is denoted by $\zeta^{a}=\{(y,\sigma,y^{\prime})\in\zeta:\sigma\in\Sigma_{o}^{a}\}$. For each transition ${tr}=(y,\sigma,y^{\prime})\in\zeta^{a}$, the attacker can change the event $\sigma$ to any string in a language $A^{M}_{\sigma}\subseteq\Sigma^{*}$, in a way similar to $A_{\sigma}$, but for transitions in $M$. Note that, since the attacker observes $\Sigma_{o}$, it knows which state $y\in Y$ it is in. Hence, when it observes an event, it knows which transition the event corresponds to. Therefore, ALTER can be implemented based on $M$.
@@ 3.3 Relation between local and global large languages
The large language generated by $(\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G$ is given by
@@
$(\mbox{by the definition of }L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G))$
@@
$(\mbox{because }\sigma\in\Sigma_{1}\setminus\Sigma_{2}\Rightarrow\sigma\not\in\Sigma\setminus\Sigma_{1}\wedge\sigma\in\Sigma\setminus\Sigma_{2})$
@@
$(\mbox{because $L_{a}(\mathcal{S}_{1}^{a}/G_{1})\parallel L_{a}(\mathcal{S}_{2}^{a}/G_{2})$ is closed })$
@@
$(\mbox{because }\sigma\in\Sigma_{1}\setminus\Sigma_{2})$
@@
$(\mbox{by Induction Hypothesis, }s\in L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G)\Rightarrow P_{2}(s)\in L_{a}(\mathcal{S}_{2}^{a}/G_{2}))$
@@
$(\mbox{by the definition of }L_{a}(\mathcal{S}_{1}^{a}/G_{1}))$
@@
$(\mbox{because }s\in L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G)\wedge s\sigma\in L(G)\Rightarrow P_{1}(s)\in L_{a}(\mathcal{S}_{1}^{a}/G_{1})\wedge P_{1}(s)\sigma\in L(G_{1}))$
@@
$(\mbox{by the definition of }L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G))$
@@
$(\mbox{because }\sigma\in\Sigma_{1}\cap\Sigma_{2}\Rightarrow\sigma\not\in\Sigma\setminus\Sigma_{1}\wedge\sigma\not\in\Sigma\setminus\Sigma_{2})$
@@
$(\mbox{because }\sigma\in\Sigma_{1}\cap\Sigma_{2})$
@@
$(\mbox{by the definition of }L_{a}(\mathcal{S}_{i}^{a}/G_{i}))$
@@
$(\mbox{because }s\in L_{a}((\mathcal{S}_{1}^{a}\wedge\mathcal{S}_{2}^{a})/G)\wedge s\sigma\in L(G)\Rightarrow P_{i}(s)\in L_{a}(\mathcal{S}_{i}^{a}/G_{i})\wedge P_{i}(s)\sigma\in L(G_{i}))$
@@ 4 Design of Coordination Control under Cyber Attacks
There exist local supervisors $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that
@@
Replace unobservable transitions in $H_{i}^{a}$ by $\varepsilon$-transitions as in Equation (5) and denote the resulting automaton as
@@
$E_{H,i}^{a}(w)=\{$ $q\in Q_{H,i}:(\exists s\in L(H_{i}))w\in\Phi_{i}^{a}(s)\wedge\delta_{H,i}(q_{i,o},s)=q\}.$
Step 3. To obtain the state estimates, we convert $H_{i,\varepsilon}^{a}$ to a deterministic automaton $H_{i,obs}^{a}$, called CA-observer, as
@@
Step 4. To ensure that the supervised system never enters illegal/unsafe states $Q_{i}\setminus Q_{H,i}$, the set of events that need to be disabled after observing $w\in\Phi_{i}^{a}(L(H_{i}))$ is given by
@@
Consider a discrete event system $G_{i},i=1,2$ under joint sensor and actuator attacks. For a nonempty closed language $P_{i}(K)\subseteq L(G_{i})$, if $P_{i}(K)$ is CA-controllable (with respect to $L(G_{i})$, $\Sigma_{uc}\cap\Sigma_{i}$, and $\Sigma^{a}_{c}\cap\Sigma_{i}$) and CA-observable (with respect to $L(G_{i})$, $\Sigma_{o}\cap\Sigma_{i}$, and $\Phi_{i}^{a}$), then the CA-supervisors $\mathcal{S}_{i,CA}$ defined in Equation (13) achieve $P_{i}(K)$, that is,